OpenStudy (anonymous):

solve for x. 2x-4y=8

OpenStudy (anonymous):

Isolate your x so add 4y to both sides to get rid of the -4y on the left side.

OpenStudy (anonymous):

2x=8-4y x=8-4y/2

OpenStudy (anonymous):

Don't give out the answer like that -_-

OpenStudy (anonymous):

if you add 4y to the right side wouldn't you have a y on that side then

OpenStudy (anonymous):

oopss sorry.. yea u have to add 4y on both sides..

OpenStudy (anonymous):

like 2x-4y+4y=8+4y on the left side the -4y+4y cancels..

OpenStudy (anonymous):

so then would you have 2x=12y?

OpenStudy (anonymous):

no 2x=8+4y now u hv to divide 2 on both sides so that u r left with x!

OpenStudy (anonymous):

u understand wht i meant?

OpenStudy (anonymous):